In a random sample of 64 people it was found that 42 of them could name all fifty states in the U.S. If two people from this sample are randomly chosen without replacement, what is the probability at least one of these two people can name all fifty states?

Round your answer to 4 decimal places.

(Hint: The complement of "at least one" is "none". Find the probability that neither of the two people chosen can name all fifty states and then use the "complements principle" to find your answer.)

please show all steps on how to figure this out.
Answer #1

Number of people who can't name all fifty states = 64 - 42 = 22

Hence,

P(At least one of these two people can name all fifty states)

= 1 - P(None of these two people can name all fifty states)

= 1 - (22/64)*(21/63)

= 0.8854